Shift Summary for Eve shift, 7/Sept/99

Miguel Barrio (mbarrio@hep.uchicago.edu)
Tue, 7 Sep 1999 00:01:03 -0500

Shift Summary for Eve shift, 7/Sept/99

On Shift: BobT, Miguel Barrio

Hours of beam available at all intensities: 7.3
Hours of beam available at nominal intensity: 7.3
Hours of E832 data (special runs not included): 6.8
Hours of special runs (laser scans, muon runs...): 0.0

Accounting of lost time:
Tape changes (circa 10 minutes).
Targeting was way off for about 20 minutes.

========================================================================
Runs ended this shift:
Run 14447 - 9.5M events - BEAM_832
Run 14448 - 13.6M events - BEAM_832 - end of tapes

Tapes brought to FCC this shift: QK*399-401, QKB140

========================================================================
Accesses during shift:
none

========================================================================
Overview of shift:
E832 data taking

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------

14447 15:56 No beam. Stopped run to check vac+-/vac00 ratio
since the HA is double counting. vac+-/vac00
ratio is ok, so trigger is ok. Problem is
likely a scaler problem.

14448 16:39 Beam's back. Started run 14448.

~17:00 VolMon2 complains about voltage in RR174/NIM1.
Readout indicates low voltage in the +6V line
(reads 5.3V).

17:07 TRD6 has problems. First, the sweep current draw
is high (~30uA). Next chamber 6 trips.

17:25 Ed calls, he tells us NOT to take a muon run
tonight (neither in the evening nor owl shifts).

18:27 VolMon2 complaining again, it now reads 5.1V.

19:10 Sure enough, VolMon2 is screaming again, now
reading 3.4V (poor thang!). Checked the NIM
crate voltages, and indeed the +6V line is low
at 5.33V. All other test points read fine.

20:15 Seed device indicates targeting is all ova da
place. We call the control room, they say they
will check it out. We stop the trigger in the
meantime.

20:27 MCR managed to get that drunk driver off of the
targetting adjustment knobs and the beam is back
on target. We start triggers again and off we
go.

21:06 While having outrageous amounts of fun checking
all the monitoring plots, we notice that RC10 is
still reading high. This is a known problem so
we don't worry about it.

21:15 Jim says we need to mount one of those hideous
QKNL tapes... QKNL33 to be precise. His voice
lacks the conviction so evident in other
audible alarms... but in any case we grant him
his wish and go mount da darn tape.

21:20 Tapes full. Run stops. We change tapes to
QK*401-404, QKB141.

14449 21:30 Started run 14449 with new cutesie tapes.

23:15 Got some Ktevana errors, and a pipeline
screw-up. BobT promptly fixes both (wow!).

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

NO MUON RUN... Woo-hoo!